OpenStudy (anonymous):

The coordinates of the midpoint of line segment AB are (-2,4). If the coordinates of point A are (7,10), what are the coordinates of point B? A) (-2 1/2,7) B) (5,6) C) (11,-2) D) (-11,-2)

OpenStudy (anonymous):

@huneya

OpenStudy (anonymous):

use the midpoint formula for this

OpenStudy (anonymous):

ok

OpenStudy (triciaal):

do you still need help and did you draw what it looks like?

OpenStudy (anonymous):

Is the answer A?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Why isn't it A?

OpenStudy (anonymous):

because a is the x is -2 is already in the midpoint coordinate

OpenStudy (triciaal):

i didn't read the question right to find the midpoint of a line segment (x1 + x2)/2 , (Y1 + Y2) / 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-2+7/2, 4+10/2 SIMPLIFIES to 5/2, 14/2

OpenStudy (anonymous):

I got 2.5 (2 1/2) + 7= 9.5

OpenStudy (anonymous):

no u dont simply

OpenStudy (anonymous):

Using the midpoint rule (x+7)/2=-2 and (y+10)/2 = 4 x = -11 y = -2 B is Pt(-11,-2)

OpenStudy (triciaal):

The coordinates of the midpoint of line segment AB are (-2,4). If the coordinates of point A are (7,10), what are the coordinates of point B? LET (x1, y1) = (7,10) let B = (x2, y2) (7 + x2)/2 = -2 (10 + y2) /2 = 4 can you finish this?

OpenStudy (anonymous):

so, the answer is D?

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

wow, that was tricky

OpenStudy (anonymous):

Thanks to all!

OpenStudy (triciaal):

not really have to pay attention to what the question is asking and know what formula to apply. example me ; if i read properly first would not ask if you already plot the coordinates

OpenStudy (triciaal):

you are welcome